OpenStudy (anonymous):

A 10−gallon container of water costs (y + 2) dollars while a 5−gallon container of water costs (y − 2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. All of y plus 2 divided by 10 equals all of y minus 2 divided by 5 What is the value, in dollars, of y?

OpenStudy (anonymous):

(y+ 2 )/ 10 =( y-2) / 5 multiply thru by LCD of 10: y + 2 = 2y - 4 y = 6 dollars

OpenStudy (anonymous):

ok with that?